# An Agentic AI Workflow to Simplify Parameter Estimation of Complex Differential Equation Systems

### v1

@@ Abstract
Parameter identification for mechanistic Ordinary Differential Equation (ODE) models underpins prediction and control in several applications, yet remains a labor-intensive and brittle process: datasets are noisy and partial, models can be stiff or misspecified, and differentiable implementations demand framework expertise. An agentic AI workflow is presented that converts a lightweight, human-readable specification into a compiled, parallel, and differentiable calibration pipeline. Users supply an XML description of the problem and fill in a Python code skeleton; the agent automatically validates consistency between spec and code, and auto-remediates common pathologies. It transforms Python callables into pure JAX functions for efficient just-in-time compilation and parallelization. The system then orchestrates a two-stage search comprising global exploration of the parameter space followed by gradient-based refinement. The result is an AD-native, reproducible workflow that lowers the barrier to advanced calibration while preserving expert control. An open-source implementation with a documented API and examples is released, enabling rapid movement from problem statement to fitted, auditable models with minimal boilerplate.
@@ 1 Introduction
A variety of strategies exist to calibrate these ODE models to data. Problem-specific heuristic methods, where techniques such as equation linearization reduce the estimation problem to a linear regression [7, 8, 9], have been demonstrated in several use cases, but suffer from robustness and generalizability issues[10]. Population-based algorithms (like genetic algorithms or swarm optimization) have been successfully used in applications such as battery modeling[11, 12], structural engineering [13], and piezoelectrics [14]. However, population-based methods suffer from poor convergence properties near local minima, and scale poorly with search space dimension. This has been addressed in the past using gradient-based search methods near local optima. In particular, derivatives obtained using automatic differentiation instead of numerical differences have been shown considerable success in solving ODE and PDE (Partial Differential Equation) model identification problems [15, 16, 17], particularly when the ODE models exhibit stiffness characteristics [10, 18]. By backpropagating through the ODE solve—via algorithmic differentiation of the integrator or adjoint sensitivity methods—one obtains exact derivatives of the time-discretized simulation [19] output with respect to parameters, compared to numerical differencing which have inherent discretization errors.
@@ 2 Proposed Workflow
Figure 1: (1(a)) The overall workflow proposed to automate ODE parameter estimation. Large parts of the overall workflow, which originally would have to be done by the user and would require expertise in programming, optimization and mathematics, can be handled by an agentic AI workflow.
@@ 2.2 Step II: Setup Error Corrections
Figure 3: (3(a)) Step II takes the completed Python functions and input XML, and generates a report containing critical errors and warnings. The agent automatically corrects the input files based on issues observed, and then reassesses the inputs. This is done in a loop until all critical errors are fixed, or a maximum number of iterations is reached.
After the user completes the Python functions, the XML and Python callables are resubmitted to the agent for cross-checking. The agent performs input validation, checking consistency between XML declarations and code (for example, checking for correct function signatures, correct array shapes within functions, and any unused parameters), detection of implementation pitfalls, and diagnosis of potentially suboptimal optimization choices (e.g., ill-scaled variables, infeasible bounds). Findings are partitioned into critical errors—conditions likely to cause solver failure—and warnings—issues that may degrade accuracy or efficiency. Based on this report, the agent automatically attempts to fix some of the identified faults, while broadcasting changes made to the user. This is done in a loop until the maximum number of agentic iterations is reached, or no more critical errors are detected. Figure 3 describes the workflow in this step; the agent reads the input XML and completed functions, and examines the setup for errors and inconsistencies. It auto-iterates to correct any critical errors, and apprises the user of more minor issues.
@@ 3.2.1 Piezoelectric Actuator Modeling
The framework is demonstrated on a piezoelectric actuator experiment[14] exhibiting rate-dependent hysteresis, modeled via the Bouc–Wen formulation and its voltage-driven variant. This benchmark couples nonlinear hysteresis with second-order mechanical dynamics, producing characteristic loops and transient behavior that the model must capture. The goal is to estimate the hysteresis and plant parameters from measured displacement under a known voltage drive.

### v2

@@ Abstract
Parameter identification for mechanistic Ordinary Differential Equation (ODE) models underpins prediction and control in several applications, yet remains a labor-intensive and brittle process: datasets are noisy and partial, models can be stiff or misspecified, and differentiable implementations demand framework expertise. An agentic AI workflow is presented that converts a lightweight, human-readable specification into a compiled, parallel, and differentiable calibration pipeline. Users supply an XML description of the problem and fill in a Python code skeleton; the agent automatically validates consistency between spec and code, and auto-remediates common pathologies. It transforms Python callables into pure JAX functions for efficient just-in-time compilation and parallelization. The system then orchestrates a two-stage search comprising global exploration of the parameter space followed by gradient-based refinement. The result is an AD-native, reproducible workflow that lowers the barrier to advanced calibration while preserving expert control. An open-source implementation with a documented API and examples is released, enabling rapid movement from problem statement to interpretable ODE models with minimal effort.
@@ 1 Introduction
A variety of strategies exist to calibrate these ODE models to data. Problem-specific heuristic methods, where techniques such as equation linearization reduce the estimation problem to a linear regression [7, 8, 9], have been demonstrated in several use cases, but suffer from robustness and generalizability issues[10]. Population-based algorithms (like genetic algorithms or swarm optimization) have been successfully used in applications such as battery modeling[11, 12], structural engineering [13], and piezoelectrics [14]. However, population-based methods suffer from poor convergence properties near local minima, and scale poorly with search space dimension. This has been addressed in the past using gradient-based search methods near local optima. In particular, derivatives obtained using automatic differentiation instead of numerical differences have been shown considerable success in solving ODE and PDE (Partial Differential Equation) model identification problems [15, 16, 17], particularly when the models exhibit stiffness characteristics [10, 18]. By backpropagating through the ODE solve—via algorithmic differentiation of the integrator or adjoint sensitivity methods—one obtains exact derivatives of the time-discretized simulation [19] output with respect to parameters, compared to numerical differencing which have inherent discretization errors.
@@ 2 Proposed Workflow
Figure 1: (1(a)) The overall workflow proposed to automate ODE parameter estimation. Large parts of the workflow, which originally would have to be done by the user and would require expertise in programming, optimization and mathematics, can be handled by agentic AI.
@@ 2.2 Step II: Setup Error Corrections
Figure 3: (3(a)) Step II takes the completed (completions highlighted) Python functions and input XML, and generates a report containing critical errors and warnings. The agent automatically corrects the input files based on issues observed, and then reassesses the inputs. This is done in a loop until all critical errors are fixed, or a maximum number of iterations is reached.
After the user completes the Python functions, the XML and Python callables are resubmitted to the agent for cross-checking. The agent performs input validation, checking consistency between XML declarations and code (for example, checking for correct function signatures, correct array shapes within functions, and any unused parameters), detection of implementation pitfalls, and diagnosis of potentially suboptimal optimization choices (e.g., ill-scaled variables, infeasible bounds). Findings are partitioned into critical errors—conditions likely to cause solver failure—and warnings—issues that may degrade accuracy or efficiency. Based on this report, the agent automatically attempts to fix some of the identified faults, while broadcasting changes made to the user. This is done in a loop until the maximum number of agentic iterations is reached, or no more critical errors are detected. Figure 3 describes the workflow in this step; the agent reads the input XML and completed functions, and examines the setup for errors and inconsistencies. It auto-iterates to correct any critical errors, and apprises the user of more minor issues. Appendix A.3 demonstrates examples of the workflow, showing how the agent corrects user inputs.
@@ 3.2.1 Piezoelectric Actuator Modeling
The framework is demonstrated on a piezoelectric actuator experiment (inspired from the study by Ha et al. [14]) exhibiting rate-dependent hysteresis, modeled via the Bouc–Wen formulation and its voltage-driven variant. This benchmark couples nonlinear hysteresis with second-order mechanical dynamics, producing characteristic loops and transient behavior that the model must capture. The goal is to estimate the hysteresis and plant parameters from measured displacement under a known voltage drive.
@@ 5 Funding Sources
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work the author(s) used GPT in order to check grammar and format language for better flow. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
@@ A.2 Constrained Neural Ordinary Differential Equations (NODE)
A.3 Examples of Agentic Corrections
This section demonstrates the benefits of the agentic workflow beyond just simplified and streamlined problem orchestration. A couple of examples are discussed where the agent identifies semantic and implementational errors in user inputs, automatically corrects them, reports the corrections and returns the corrected files.
The correction step itself is divided into two parts, report generation and error correction
A.3.1 Robertson’s Equations
The agent is given the user input files corresponding to the Robertson system, and asked to generate a report on their correctness. The agent identifies two semantic errors, automatically rewrites the files with corrections, and then re-checks the files. Once it finds no critical errors, it ends step II of the workflow, allowing the user to proceed to step III. Figure 10 shows the workflow of agentic correction of input files based on suboptimal inputs. This correction loop runs for only 1 iteration, however it can be configured to be run for more iterations if necessary.
Figure 10: Workflow of correction of suboptimal inputs in the Robertson parameter estimation problem, in step II of the overall workflow. The agent examines the input files and corrects errors it considers significant, and edits the input files while apprising the user of the changes made. It then again reads the input files, and saves a final report.
A.3.2 Battery Thermal Runaway Modeling
The agent is given the user input files corresponding to the thermal runaway system, and asked to generate a report on their correctness. The agent identifies missing information in the input XML and an incorrectly named function. The agent corrects them before re-checking the inputs, and given it sees no more critical errors, ends step II. The user may choose to address the generated warnings. Figure 12 shows the workflow of agentic correction of input files based on suboptimal inputs. This correction loop runs for only 1 iteration, however it can be configured to be run for more iterations if necessary.
Figure 12: Workflow of correction of suboptimal inputs in the battery thermal runaway model parameter estimation problem, in step II of the overall workflow.